# Continuous-time multifarious systems - Part II: non-reciprocal multifarious self-organization

### v2

@@ I Introduction
The natural world is filled with intricate structures that emerge from the self-assembly of simpler building blocks, ranging from the folding of proteins to the formation of cellular membranes. Harnessing these processes for synthetic materials has been a longstanding goal[1, 2], as it opens avenues for creating systems capable of dynamic adaptation. For instance, multifarious self-assembly systems can be designed to assemble multiple distinct target structures from a shared pool of building blocks [3, 4, 5, 6, 7, 8]. This is an important concept in natural self-assembly. For example, the myriad of assembled proteins are all formed from the same set of around twenty amino acids. In recent years, research into self-organizing systems has revealed that incorporating mechanisms for disassembly and reassembly can further enhance the complexity and functionality of these materials. Many mechanisms for dynamic self-organization have been explored[9, 10, 11, 12], including allosteric interactions [13, 14, 15, 16] and modifying the chemical potential of tiles [17, 18, 19]. One promising approach in this domain involves non-reciprocal interactions, which break traditional symmetry in interactions and allow for more versatile self-organization [20, 21, 22, 23, 24, 25, 26, 27, 28, 5, 29, 30].
@@ II Model
The reciprocal interactions are constructed such that the system is able to assemble the different structures encoded in the system given a suitable seed. To do this, each tile species interacts reciprocally with strength $\varepsilon$ to any tile species it neighbors in a structure[54, 7, 5]. We denote the set of all specific interactions imposed by the $m$ structures $S^{(1)},S^{(2)},\dots,S^{(m)}$ stored in the system as ${\cal I}^{\mathrm{r}}$. This is the union of smaller sets, each containing the specific interactions imposed by one of the target structures,
@@
[…]
@@
Figure 2: All the (a) reciprocal and (b) non-reciprocal interactions involving tile species 1 for an example system storing two $4\times 4$ structures with a shifting sequence from Structure 1 to Structure 2.
@@
For the non-reciprocal interactions, we use the same procedure as in previous work on non-reciprocal multifarious self-organization [5, 7]. First, for a given system we impose a shifting sequence $S^{1}\to S^{2}\to\dots\to S^{q}$. The non-reciprocal interactions are then designed such that the system is able to shift between these structures in the given sequence. To do this, tiles in a given structure interact non-reciprocally with tiles in the previous structure in the sequence. In particular, they have a non-reciprocal interaction of strength $\lambda$ to tiles they would neighbor if the structure was overlaid onto the preceding structure in the shifting sequence. An illustration of the non-reciprocal interactions is shown in Fig. 2(b). For a more precise definition of the non-reciprocal interactions, consider a single shift $S^{\ell}\rightarrow S^{\ell+1}$. A tile will approach and interact with the neighbors of its counterpart in the previous pattern in the sequence. The set of all non-reciprocal interactions imposed for this shift is
@@
$S^{\ell+1}_{i,j}\nearrow S^{\ell}_{i+1,j},\;S^{\ell+1}_{i,j}\searrow S^{\ell}_{i,j-1},\;S^{\ell}_{i,j+1}\nwarrow S^{\ell+1}_{i,j}\;\Big{\}}.$ (4)
The set of all non-reciprocal interactions for the full sequence is denoted ${\cal I}^{\mathrm{nr}}\equiv I^{\mathrm{nr}}(S^{1}\rightarrow S^{2})\cup\dots\cup I^{\mathrm{nr}}(S^{q-1}\rightarrow S^{q})$. We define
@@
[…]
@@
We assume that the reservoir is very large, and set the chemical potential of each tile species to the same constant value $\mu$. By definition the chemical potential of the solvent is 0. We use a discrete-time Monte Carlo algorithm and a continuous-time Gillespie algorithm to simulate the time evolution of these models. The details of the simulation methods are presented in Appendix A.
Figure 4: Non-equilibrium state diagram using continuous-time simulations, with $\lambda=10$ and shifting sequence $S^{1}\to S^{2}\to S^{3}$. For each value of $(-\mu,\varepsilon)$, a system with $m=3$ stored $40\times 40$ structures is simulated in the $80\times 80$ lattice with $10^{8}$ reaction steps. Left: Error, density, and energy at the end of the simulations. The different state regions are marked, with MA standing for multifarious assembly, and SS for shape-shifting. The shape-shifting region is also marked by the dashed red lines. Right: Snapshots at the end of the simulations, colored by structure as described in Appendix B.
@@ III Non-equilibrium multifarious self-organization
The shape-shifting state is characterized by a low error, with the overlaps for each structure rising and falling along the pre-defined shifting sequence. The dynamics of the overlaps is shown in Fig. 3, alongside snapshots of shape-shifting. In the multifarious-self assembly state, the system can successfully assemble structures (characterized by the low error), but if the non-reciprocal interaction strength $\lambda$ is too low, there will be no shape-shifting.
The chimera state occurs when the reciprocal interaction strength $\varepsilon$ is too high and can be identified by high error, high density, and low energy. The system builds uncontrollably, building unwanted structures attached to the desired ones. A sub-state of the chimera regime is homogeneous nucleation, where large domain sizes lead to a particularly low energy. The liquid and dispersion states are characterized by a high error and energy, and can be distinguished via the density. The liquid state has a high density, whilst the dispersion state has a low density.
In Fig. 4, we show the state diagram obtained using continuous-time simulations in non-equilibrium systems, with $\lambda=10$. The different states described are marked in the corresponding locations on the state diagram. The boundary line for instability towards chimeric structures is the same as that found in equilibrium systems[54]. In particular, this boundary is different compared with discrete-time simulations, due to the faster chimeric growth dynamics in the continuous-time simulations[54]. As found previously in discrete-time simulations[5], increasing the non-reciprocal interaction strength pushes the liquid-dispersion transition point to larger $-\mu$ and the liquid-chimera transition point to larger $\varepsilon$.
@@ IV.1 Nucleation
$t_{N}\left({\frac{2}{3}\varepsilon<\lambda<\varepsilon}\right)=\frac{1}{L^{2}}e^{-\mu-7\lambda+4\varepsilon}+e^{-\mu-4\lambda+2\varepsilon},$ (8)
is plotted in Fig. 5(b) and agrees well with nucleation times measured in simulations. We find a crossover in scaling from $t_{N}\sim e^{-7\lambda}$ for small $\lambda$ to $t_{N}\sim e^{-4\lambda}$ for large $\lambda$. The first term in the right-hand side of Eq. (8) is the contribution coming from the time to create new nucleation seeds. For small $\lambda$ this is slow, and so this term dominates. As $\lambda$ increases, the time to create nucleation seeds drops, leading to the observed crossover in scaling.
@@
where $f<0.5$ is the target fraction of $S^{2}$ which we set to $f=0.2$. The derivation of this expression is given in Appendix D.1. In discrete-time simulations this parameter range is always unstable[5], destroying any structure in the system. In continuous-time simulations some systems are able to successfully shape-shift using this mechanism. However, even in continuous-time simulations it does not give stable intermediate structures, and can also completely destabilize systems, for example with cyclic shifting sequences.
@@ IV.2 Interface growth
To measure interface growth we once again consider a system with two stored structures ($S^{1}$ and $S^{2}$) and a shifting sequence $S^{1}\to S^{2}$. We initialize the system with a flat interface by placing a slab of $S^{2}$ underneath a slab of $S^{1}$. We use hard-wall and periodic boundary conditions perpendicular and parallel to the interface respectively. In order to measure the interface over long times we use interface displacement [55, 7], which overcomes the issue of a finite system size. Furthermore, since we are interested in characterizing the behavior of a single interface, when measuring interface velocity we only update tiles within a certain range of the interface to prevent spurious nucleation and the formation of new interfaces. On the left-hand side of Fig. 6(a) we show interface growth in a system with $\lambda<\varepsilon$. In this case the interface moves steadily upwards as the $S^{2}$ region grows. On the right-hand side of Fig. 6(a) we show an example of the system evolution when $\lambda>\varepsilon$. As discussed previously, in this case $S^{2}$ tiles arrive across the whole lattice, independently of the already present $S^{2}$ region, and we do not see interface growth.
@@ IV.3 Discussion
In discrete-time simulations the timescale of nucleation and interface growth is generally constant with respect to $\lambda$ [7]. This is because the process of proposing a favorable reaction, which involves picking both the correct position and the correct new tile, is a very slow process for moderate to large system sizes. Therefore, the timescales are not strongly affected by $\lambda$, which only modifies the strength of an already quite high selection probability. In contrast, in continuous-time simulations a reaction is always selected, with the most favorable reactions being the most likely reactions to be picked. Furthermore, the time between reactions depends on the total reaction rate, which depends on $\lambda$. Therefore, in continuous-time simulations we see very strong dependencies of the timescales on $\lambda$.
@@ V Conclusions
Extending the model to different types of interactions and environment conditions could broaden its applicability to real-world scenarios, such as adaptive materials or programmable matter. Aside from creating dynamic states, non-equilibrium strategies have also been exploited to make systems more robust or efficient than their equilibrium counterparts [56, 57, 58], including the application of non-reciprocal interactions [7]. Overall, our findings contribute to the growing field of non-equilibrium self-assembly, offering a pathway to design systems capable of dynamic structural reconfiguration.
@@ A.1 Discrete-time simulations
$p_{\mathrm{acc}}=\min[1,e^{\Lambda-\Delta\mathcal{H}}].$ (11)
Here the equilibrium total energy
$\mathcal{H}=\sum_{\langle\alpha,\beta\rangle}U^{\rm r}_{\sigma_{\alpha}\square\sigma_{\beta}}-\mu\sum_{\alpha}(1-\delta_{0,\sigma_{\alpha}}),$ (12)
has contributions from the reciprocal bonds and chemical potentials. The non-equilibrium drive
$\Lambda=R_{\sigma_{i-1,j}\swarrow\sigma^{\prime}}^{\mathrm{nr}}+R_{\sigma^{\prime}\nearrow\sigma_{i+1,j}}^{\mathrm{nr}}+R_{\sigma^{\prime}\searrow\sigma_{i,j-1}}^{\mathrm{nr}}+R_{\sigma_{i,j+1}\nwarrow\sigma^{\prime}}^{\mathrm{nr}},$ (13)
contains the contribution from the non-reciprocal interactions. If the reaction is accepted then the system is correspondingly updated, otherwise the system remains unchanged. This processes is repeated for the desired simulation time.
@@ A.2 Continuous-time simulation
Prob(selecting $\text{Prob}(\text{selecting }r)=k_{r}/k_{T},$ (17)
@@ Appendix C Order parameters
To find the overlaps, we first find the largest connected cluster of tiles $G$. For each structure we then calculate $A_{\ell}=G\cup S^{\ell}$, where $S^{\ell}$ is $\ell$th structure located at the center of the lattice. The overlap for each structure is then given by
@@ Appendix D Analytical timescale calculations
Our calculation technique relies on several simplifying assumptions which make the calculation tractable by hand. Firstly we focus only on structural identity, without needing to identify each particular tile species. This essentially neglects the chance that a tile shares neighbors in multiple structures, which becomes vanishingly small in the limit of large $l$. We also neglect transitions involving empty sites, assuming that the chemical potential is sufficiently low as compared to the interactions that transitions involving empty sites will be subdominant. Typically a few transitions will have by far the largest rate, such that the average time for that transition is approximately the inverse of the sum of the rates of these dominant transitions. These assumptions greatly simplify the calculations, as a comprehensive enumeration might not be tractable. Despite these assumptions, the resulting calculations provide a good approximation. The main reason for this is the exponentiation in the rates. This means that particular transitions quickly become very dominant, which can be used to simplify rate calculations as described above. Furthermore, for each process there is typically one rate-limiting step (RLS) which dominates the timescales. In interface growth the RLS is initiating a new layer. In nucleation the RLS is forming a critical seed. This means that the scaling in the timescales is typically due to one key process, and once this process has been considered in the calculation, the correct scaling will emerge.

### v3

@@ I Introduction
The natural world is filled with intricate structures that emerge from the self-assembly of simpler building blocks, ranging from the folding of proteins to the formation of cellular membranes. Harnessing these processes for synthetic materials has been a longstanding goal[1, 2], as it opens avenues for creating systems capable of dynamic adaptation. For instance, multifarious self-assembly systems can be designed to assemble multiple distinct target structures from a shared pool of building blocks [3, 4, 5, 6, 7, 8]. This is an important concept in natural self-assembly. For example, the myriad of assembled proteins are all formed from the same set of around twenty amino acids. In recent years, research into self-organizing systems has revealed that incorporating mechanisms for disassembly and reassembly can further enhance the complexity and functionality of these materials. Many mechanisms for dynamic self-organization have been explored[9, 10, 11, 12, 13, 14], including allosteric interactions [15, 16, 17, 18] and modifying the chemical potential of tiles [19, 20, 21]. One promising approach in this domain involves non-reciprocal interactions, which break traditional symmetry in interactions and allow for more versatile self-organization [22, 23, 24, 25, 26, 27, 28, 29, 30, 5, 31, 32].
@@
Figure 2: All the (a) reciprocal and (b) non-reciprocal interactions involving tile species 1 for an example system storing two $4\times 4$ structures with a shifting sequence from Structure 1 to Structure 2. The green rectangular markers show the reciprocal interactions, and the red triangular markers show the non-reciprocal interactions.
@@ II Model
Multifarious systems are designed to assemble multiple different structures using a shared set of components. In the equilibrium multifarious self-assembly model [3], structures are encoded within the interactions between building blocks - a molecular analogy of a Hopfield network [56]. A structure can be ‘retrieved’ using an appropriate seed, upon which the system assembles the full corresponding structure.
Including non-reciprocal interactions between the building blocks, where the interactions are not symmetric, breaks detailed balance. This leads to a fundamentally non-equilibrium model. The resulting non-reciprocal multifarious self-organization model exhibits a variety of non-trivial behaviors. For example, systems can be designed to shape-shift, with assembled structures autonomously reorganizing into new structures, following a custom predefined sequence of structures. The non-reciprocal interactions can also be used to drive systems out of otherwise extremely long-lived kinetic traps [7].
@@
The reciprocal interactions represent bonds that form between neighboring tiles, lowering the system energy. They are constructed such that the system is able to assemble the different structures encoded in the system given a suitable seed. To do this, each tile species interacts reciprocally with strength $\varepsilon$ to any tile species it neighbors in a structure[57, 7, 5]. We denote the set of all specific interactions imposed by the $m$ structures $S^{(1)},S^{(2)},\dots,S^{(m)}$ stored in the system as ${\cal I}^{\mathrm{r}}$. This is the union of smaller sets, each containing the specific interactions imposed by one of the target structures,
@@
Here $\alpha$ and $\beta$ are lattice coordinates running over nearest neighbors. A specific interaction belongs to the set ${\cal I}^{\mathrm{r}}$ if at least one target structure favors that interaction. To enable clear notation we rotate the lattice as shown in Fig. 2(a). $\square\in\{\diagdown,\diagup\}$ are the vertical and horizontal directions of the specific interactions. Putting everything together, the interaction matrix is then
@@
The reciprocal interactions and the chemical potential $\mu$, which is defined later, form the equilibrium total energy
$\mathcal{H}=\sum_{\langle\alpha,\beta\rangle}U^{\rm r}_{\sigma_{\alpha}\square\sigma_{\beta}}-\mu\sum_{\alpha}(1-\delta_{0,\sigma_{\alpha}}).$ (4)
Using this energy in discrete-time Monte Carlo or continuous-time Gillespie simulations would model a system in equilibrium. We now describe how the non-reciprocal interactions are constructed, and the resulting non-equilibrium drive.
@@
For the non-reciprocal interactions, we use the same procedure as in previous work on non-reciprocal multifarious self-organization [5, 7]. The non-reciprocal interactions represent biases in the Monte Carlo acceptance probability and in the Gillespie rates that break detailed balance. First, for a given system we impose a shifting sequence $S^{1}\to S^{2}\to\dots\to S^{q}$. The non-reciprocal interactions are then designed such that the system is able to shift between these structures in the given sequence. To do this, tiles in a given structure interact non-reciprocally with tiles in the previous structure in the sequence. In particular, they have a non-reciprocal interaction of strength $\lambda$ to tiles they would neighbor if the structure was overlaid onto the preceding structure in the shifting sequence. An illustration of the non-reciprocal interactions is shown in Fig. 2(b). For a more precise definition of the non-reciprocal interactions, consider a single shift $S^{\ell}\rightarrow S^{\ell+1}$. A tile will approach and interact with the neighbors of its counterpart in the previous pattern in the sequence. The set of all non-reciprocal interactions imposed for this shift is
@@
$S^{\ell+1}_{i,j}\nearrow S^{\ell}_{i+1,j},\;S^{\ell+1}_{i,j}\searrow S^{\ell}_{i,j-1},\;S^{\ell}_{i,j+1}\nwarrow S^{\ell+1}_{i,j}\;\Big\},$ (5)
where $(i,j)$ are the indices of the two-dimensional lattice coordinates. The set of all non-reciprocal interactions for the full sequence is denoted ${\cal I}^{\mathrm{nr}}\equiv I^{\mathrm{nr}}(S^{1}\rightarrow S^{2})\cup\dots\cup I^{\mathrm{nr}}(S^{q-1}\rightarrow S^{q})$. We define
@@
where $\blacksquare\in\{\searrow,\nwarrow,\nearrow,\swarrow\}$ represents all possible specific non-reciprocal interaction directions. The non-reciprocal interactions give rise to a non-equilibrium drive. For the reaction changing tile $\sigma_{i,j}$ to tile $\sigma^{\prime}$, at lattice site $(i,j)$, the drive is
$\Lambda=R_{\sigma_{i-1,j}\swarrow\sigma^{\prime}}^{\mathrm{nr}}+R_{\sigma^{\prime}\nearrow\sigma_{i+1,j}}^{\mathrm{nr}}+R_{\sigma^{\prime}\searrow\sigma_{i,j-1}}^{\mathrm{nr}}+R_{\sigma_{i,j+1}\nwarrow\sigma^{\prime}}^{\mathrm{nr}}.$ (7)
As explained further in Appendix A, the non-equilibrium drive is used in the Monte Carlo acceptance probability and in the Gillespie rates, breaking detailed balance and resulting in a fundamentally non-equilibrium model.
@@
We assume that the reservoir is very large, and set the chemical potential of each tile species to the same constant value $\mu$. By definition the chemical potential of the solvent is 0.
We use a discrete-time Monte Carlo algorithm and a continuous-time Gillespie algorithm to simulate the time evolution of these models. The details of the simulation methods are presented in Appendix A.
Figure 4: Non-equilibrium state diagram using continuous-time simulations, with $\lambda=10$ and shifting sequence $S^{1}\to S^{2}\to S^{3}$. For each value of $(-\mu,\varepsilon)$, a system with $m=3$ stored $40\times 40$ structures is simulated in the $80\times 80$ lattice with $10^{8}$ reaction steps. Left: Error, density, and energy at the end of the simulations. The different state regions are marked, with MA standing for multifarious assembly, and $S$ for shape-shifting. The shape-shifting region is also marked by the dashed red lines. Right: Snapshots at the end of the simulations, colored by structure as described in Appendix B.
@@ III Non-equilibrium multifarious self-organization
The shape-shifting state is characterized by a low error, with the overlaps for each structure rising and falling along the pre-defined shifting sequence. The dynamics of the overlaps is shown in Fig. 3, alongside snapshots of shape-shifting. In the multifarious self-assembly state, the system can successfully assemble structures (characterized by the low error), but if the non-reciprocal interaction strength $\lambda$ is too low, we do not see shape-shifting. The chimera state occurs when the reciprocal interaction strength $\varepsilon$ is too high and can be identified by high error, high density, and low energy. The system builds uncontrollably, building unwanted structures attached to the desired ones. The liquid and dispersion states are characterized by a high error and energy, and can be distinguished via the density. The liquid state has a high density, whilst the dispersion state has a low density.
In Fig. 4, we show the state diagram obtained using continuous-time simulations in non-equilibrium systems, with $\lambda=10$. The different states described are marked in the corresponding locations on the state diagram. The boundary line for instability towards chimeric structures is the same as that found in equilibrium systems[57]. In particular, this boundary is different compared with discrete-time simulations, due to the faster chimeric growth dynamics in the continuous-time simulations[57]. As found previously in discrete-time simulations[5], increasing the non-reciprocal interaction strength pushes the liquid-dispersion transition point to larger $\mu$ and the liquid-chimera transition point to larger $\varepsilon$.
@@ IV.1 Nucleation
$t_{N}\left({\frac{2}{3}\varepsilon<\lambda<\varepsilon}\right)=\frac{1}{L^{2}}e^{-\mu-7\lambda+5\varepsilon}+e^{-\mu-3\lambda+\varepsilon},$ (10)
is plotted in Fig. 5(b) and agrees well with nucleation times measured in simulations. The first term in the right-hand side of Eq. (10) is the contribution coming from the time to create new nucleation seeds. For small $\lambda$ this is slow, and so this term dominates. This results in the observed $t_{N}\sim e^{-7\lambda}$ scaling. As $\lambda$ increases, the time to create nucleation seeds drops, and the second term becomes relevant. As $\lambda$ increases beyond $\varepsilon$, which we discuss in the following paragraph, we see a crossover to $t_{N}\sim e^{-4\lambda}$ scaling.
@@
where $f<0.5$ is the target fraction of $S^{2}$ which we set to $f=0.2$. The derivation of this expression is given in Appendix D.1. In this parameter range shape-shifting is generally not stable, as discussed further in Appendix D.1.
@@ IV.2 Interface growth
To measure interface growth we once again consider a system with two stored structures ($S^{1}$ and $S^{2}$) and a shifting sequence $S^{1}\to S^{2}$. We initialize the system with a flat interface by placing a slab of $S^{2}$ underneath a slab of $S^{1}$. We use hard-wall and periodic boundary conditions perpendicular and parallel to the interface respectively. In order to measure the interface over long times we use interface displacement [58, 7], which overcomes the issue of a finite system size. Furthermore, since we are interested in characterizing the behavior of a single interface, when measuring the interface velocity we do not update any tiles further than five vertical lattice sites above the highest point of the interface or five below the lowest point of the interface. This prevents spurious nucleation and the formation of new interfaces [7]. On the left-hand side of Fig. 6(a) we show interface growth in a system with $\lambda<\varepsilon$. In this case the interface moves steadily upwards as the $S^{2}$ region grows. On the right-hand side of Fig. 6(a) we show an example of the system evolution when $\lambda>\varepsilon$. As discussed previously, in this case $S^{2}$ tiles arrive across the whole lattice, independently of the already present $S^{2}$ region, and we do not see interface growth.
@@ IV.3 Discussion
In discrete-time simulations the timescale of nucleation and interface growth is generally constant with respect to $\lambda$.[7] This is because the process of proposing a favorable reaction, which involves picking both the correct position and the correct new tile, is a very slow process for moderate to large system sizes. Therefore, the timescales are not strongly affected by $\lambda$, which only modifies the strength of an already quite high selection probability. In contrast, in continuous-time simulations a reaction is always executed, with the most favorable reactions being the most likely reactions to be picked. Furthermore, the time between reactions depends on the total reaction rate, which depends on $\lambda$. Therefore, in continuous-time simulations we see very strong dependencies of the timescales on $\lambda$. Whilst some systems show similar dynamics between different simulation methods [59, 60, 61], in the system studied here, over the timescales of assembly and shape-shifting we find that the kinetics are not identical between the discrete-time Monte Carlo and the continuous-time Gillespie simulations. As explained, a key contributor of this is the fact that we are dealing with a multi-component system, with a large number of components. This is true even when detailed balance is obeyed ($\lambda=0$), as explored in the previous work of this series [57]. Nonetheless the same key behaviors, including in the non-equilibrium system, are observed using either of the two different simulation methods.
@@ V Conclusions
Interestingly, we find that the continuous-time Gillespie algorithm leads to different kinetics compared to the discrete-time Monte Carlo algorithm. For example, in the continuous-time simulations the timescales of nucleation and shape-shifting depend strongly on the non-reciprocal interaction strength $\lambda$. In discrete-time simulations the dependence of timescales on $\lambda$ is typically piecewise flat [5, 7]. Despite these differences, the same fundamental behaviors are observed using either simulation technique.
Extending the model to different types of interactions and environment conditions could broaden its applicability to real-world scenarios, such as adaptive materials or programmable matter. Aside from creating dynamic states, non-equilibrium strategies have also been exploited to make systems more robust or efficient than their equilibrium counterparts [62, 63, 64], including the application of non-reciprocal interactions [7]. A relevant experimental system to our work is that of chemically active phoretic colloids [22, 23, 24, 27, 28, 29, 30, 65, 66, 67]. These colloids produce or consume chemicals, and respond to gradients in the resulting chemical fields[68]. Mixing colloids with different chemical activities and mobilities generically leads to non-reciprocal interactions. Highly interesting self-organization behaviors have been demonstrated in colloidal systems, both experimentally and in theory. It would be interesting to explore how the dynamics we investigate in this work could be applied to such systems. Overall, our findings contribute to the growing field of non-equilibrium self-assembly, offering a pathway to design systems capable of dynamic structural reconfiguration.
@@ A.1 Discrete-time simulations
$p_{\mathrm{acc}}=\min[1,e^{\Lambda-\Delta\mathcal{H}}],$ (13)
where $\mathcal{H}$ is defined in Eq. (4) and $\Lambda$ is defined in Eq. (7). If the reaction is accepted then the system is correspondingly updated, otherwise the system remains unchanged. This processes is repeated for the desired simulation time.
@@ A.2 Continuous-time simulation
$\text{Prob}(\text{selecting }r)=k_{r}/k_{T},$ (17)
@@ Appendix C Order parameters
[…]
@@ Appendix D Analytical timescale calculations
Our calculation technique relies on several simplifying assumptions which make the calculation tractable by hand. Firstly we focus only on structural identity, without needing to identify each particular tile species. This essentially neglects the chance that a tile shares neighbors in multiple structures, which becomes vanishingly small in the limit of large $l$. We also neglect transitions involving empty sites, assuming that these transitions are subdominant. This can be seen, for example, in the snapshots shown in Fig. 5(a) and Fig. 6(a), in which there are no empty lattice sites. Typically a few key transitions will have by far the largest rate, such that the average time for that transition is approximately the inverse of the sum of the rates of these dominant transitions. For different parameter values, different transitions dominate. Based on this, we divide the calculations of timescales into different parameter ranges. These assumptions greatly simplify the calculations, as a comprehensive enumeration might not be tractable. Despite these assumptions, the resulting calculations provide a good approximation. The main reason for this is the exponentiation in the rates. This means that particular transitions quickly become very dominant, which can be used to simplify rate calculations. Furthermore, for each process there is typically one rate-limiting step (RLS) which dominates the timescales [72]. In nucleation the RLS is forming a critical seed (reactions $\mathcal{A}_{N}\to\mathcal{B}_{N}\to\mathcal{B}_{N}\to\mathcal{C}_{N}$ in Fig. 9). In interface growth the RLS is initiating a new layer (reaction $\mathcal{A}_{I}$ in Fig. 10). This means that the scaling in the timescales is typically due to one key process, and once this process has been considered in the calculation, the correct scaling will emerge.